OpenStudy (anonymous):

In an experiment using 30 mice, the sample proportion of the mice that gained weight after a drug injection is 0.65. What is the 99.7% confidence interval for the actual proportion in the population? A. between 0.602 and 0.697 B. between 0.563 and 0.737 C. between 0.476 and 0.824 D. between 0.389 and 0.911

OpenStudy (bigdaddyharsha):

According to your multiple choice answers you have to select from, the answer is between 0.563 and 0.737. Thank you for asking.

jimthompson5910 (jim_thompson5910):

Confidence Interval for a Proportion \[\Large L = \hat{p} + z_c*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\] \[\Large U = \hat{p} - z_c*\sqrt{\frac{\hat{p}*(1-\hat{p})}{n}}\] (L,U) is the confidence interval L being the lower bound, U being the upper bound \(\Large \hat{p}\) (read as "p-hat") is the sample proportion. In this case, \(\Large \hat{p} = 0.65\) n = 30 is the sample size \(\Large z_c\) is the critical z value (found using a calculator or a table)